OpenStudy (anonymous):

If ln(a)=2, ln(b)=3, and ln(c)=5, evaluate the following: ln [(a^3)/(b^-1)(c^2)]

OpenStudy (unklerhaukus):

\[\log(xy/z)=\log x+\log y -\log z\]